OpenStudy (anonymous):

The rearrangement of methyl isonitrile (CH3NC) to acetonitrile (CH3CN) is a first-order reaction and has a rate constant of 5.11×10−5s−1 at 472 K. If the initial concentration of CH3NC is 3.50×10−2M . Part A. What is the molarity of CH3NC after 1.80h ? Part B. How many minutes does it take for the CH3NC concentration to drop to 2.90×10−2M ? Part C. How many minutes does it take for 20% of the CH3NC to react?

OpenStudy (anonymous):

I have no idea how to approach this problem. I know that Molarity= mol/L for part A, but then I don't know how to convert it to hours from there.

OpenStudy (abb0t):

start by changing 1.8 hours to seconds. Your rate constant is in seconds, so you want to keep this consistent.

OpenStudy (abb0t):

\(1^{st}\) order kinetics: \(k = \frac{1}{t}ln(\frac{a}{a_0})\) solve for \(a_0\) where, \(a_0\)

OpenStudy (anonymous):

ok so 1.8 hours is 6480 seconds. is that my a_o?

OpenStudy (anonymous):

itll be (5.11*10^-5)*(472)=ln(a/6480)....I get confused on what to do with the ln part to solve for a.